# Gaussian Mapping for Evolving Scenes

### v1

@@ Abstract
Mapping systems with novel view synthesis (NVS) capabilities are widely used in computer vision, with augmented reality, robotics, and autonomous driving applications. Most notably, 3D Gaussian Splatting-based systems show high NVS performance; however, many current approaches are limited to static scenes. While recent works have started addressing short-term dynamics (motion within the view of the camera), long-term dynamics (the scene evolving through changes out of view) remain less explored. To overcome this limitation, we introduce a dynamic scene adaptation mechanism that continuously updates the 3D representation to reflect the latest changes. In addition, since maintaining geometric and semantic consistency remains challenging due to stale observations disrupting the reconstruction process, we propose a novel keyframe management mechanism that discards outdated observations while preserving as much information as possible. We evaluate Gaussian Mapping for Evolving Scenes (GaME) on both synthetic and real-world datasets and find it to be more accurate than the state of the art.
@@ 1 Introduction
Visual mapping enables a machine to build a 3D representation of its surroundings based on its camera input. Imagine an autonomous car navigating a busy city or a robot exploring a cluttered room. For these systems to move safely and make decisions, they need a reliable understanding of their surroundings. Over the years, visual mapping systems have advanced [24, 18], learning to handle complex scenes with increasing accuracy. This capability forms the backbone for various systems from self-driving cars to virtual reality glasses, where spatial understanding is crucial for navigation, planning, and interaction.
@@
Recent mapping systems have been enhanced with novel view synthesis (NVS) capabilities [18, 11], allowing them to generate realistic, immersive views of scenes. This allows more detailed scene exploration and supports the creation of high-quality virtual environments. These approaches typically assume that the scene is static such that optimization over multiple frames is well conditioned. However, real-world environments are rarely static, but include both short-term dynamic effects (i.e. things moving within view of the camera) as well as long-term dynamics (i.e., the scene evolving through changes outside the view of the camera). While some of the most recent NVS approaches address short-term dynamic objects [39, 33], the latter remains less explored. As a result, modern reconstruction methods with NVS capabilities struggle to capture changes as the scene evolves over time, preventing them from being deployed in long-term reconstruction pipelines.
@@
During long-term mapping, systems must continuously update the 3D representation to capture the scene’s evolution and maintain reliable operation. This is especially challenging compared to classical multi-session mapping as changes can occur at any time during data collection. Recently, initial methods have started to address this challenge. A first approach is presented in Panoptic Multi-TSDFs [26], building semantically consistent submaps and reasoning about changes on the level of submaps. The following works [6, 21, 27] similarly focus on object-level mapping to handle evolving scenes. However, all these methods rely on map representations that cannot easily be re-rendered, preventing their use in AR/VR or digital maps where realistic rendering is essential.
This work addresses the challenge of long-term NVS mapping using 3D Gaussian Splatting (3DGS) in evolving scenes. The key insight of our approach is that environment changes are not random but typically follow semantically consistent patterns. We therefore integrate semantic consistency with the inherent properties of 3DGS to efficiently detect and adapt to environment changes in the incrementally built 3DGS model. As multi-view optimization is essential for accurate 3DGS mapping, we introduce a keyframe management method that appropriately masks stale areas in keyframes to retain useful information while accounting for changes, resulting in a well-conditioned 3DGS optimization process even in evolving scenes. We make the following contributions:
@@
  • We present GaME, the first NVS-capable mapping system for long-term evolving scenes.
@@
  • An efficient keyframe management strategy for accurate 3D reconstruction through changes.
  • We thoroughly evaluate GaME on synthetic and real-world data, showing a performance increase of 90+in depth and 20+in color rendering. We release the code open-source [^1] Released upon acceptance..
@@ 2 Related Work
3D Change Detection. The goal of change detection is to handle long-term dynamic effects, i.e., changes to the scene occurring outside the view of the sensor. Typically, this is addressed in a multi-session setting. Most prominently, once a map for each session is constructed, changes can be identified through geometric scene differencing [1, 5, 13, 25]. To achieve an object-level change understanding, this has been extended using semantic information [14], where recent trends focus on the extraction of more specialized object features, including learned shape descriptors [29, 7, 25], neural object representations [6, 40], and language embeddings [23]. Most related to us, Lu et al. [16] recently presented a 3D Gaussian Splatting-based [12] approach by re-rendering the scene to newly collected views and using EfficientSAM [32] for 2D change detection. However, the assumption that the scene is static during each session and offline processing is highly limiting, as changes in human-centric and evolving scenes can occur at any time during the mapping process. Conversely, our mapping is designed to operate online and does not impose any limitations on the scene changes.
@@
Online Long-Term Reconstruction. Recently, initial methods have started to address this more general online problem. A first approach is presented by Schmid et al. [26], which generates locally and semantically consistent submaps and incrementally reasons about changes at the submap level. Fu et al. [6] propose neural object descriptors to build an object-level pose graph and detect changes in the graph configuration. This has recently been extended with $SE(3)$-equivariant descriptors [8]. Qian et al. [21] present a frame-to-map-tracking approach, using a probabilistic update rule to detect object-level changes, which is further extended to a variational factor-graph approach in [22]. A unified short- and long-term dynamic reconstruction formulation is presented in Khronos [27], where objects are reconstructed locally and verified for changes using a library of rays. Nonetheless, these methods rely on map representations that cannot easily be re-rendered. In contrast, GaME provides scene reconstruction capable of real-time color and depth rendering.
@@
Dynamic Gaussian Splatting. 3D Gaussian Splatting  [12] has revolutionized novel view synthesis (NVS) by enabling photorealistic, real-time rendering at over 100 FPS. Compared to neural radiance fields [19], 3DGS is significantly more memory-efficient and faster to optimize. The problem of dynamic or 4D GS has attracted wide interest. A series of works [3, 17, 31, 36] optimize a canonical set of Gaussians from the initial frame and model temporal variations through a deformation field. However, these methods are limited to short video sequences, as they cannot introduce new Gaussians after the initial frame. Another class of approaches [4, 10, 35] directly models temporal Gaussians that can exist over subsets of frames. Despite their improved flexibility, these methods require offline optimization and multi-view input, making scalability a significant bottleneck for high-quality dynamic reconstruction. In contrast, our mapping operates online using only a single RGB-D camera.
@@
Online Gaussian Mapping. Most related to us, 3D Gaussian Splatting has sparked a wave of online RGB-D mapping methods [18, 11, 34, 9, 37] that adopt 3D Gaussians as their primary scene representation. While these methods perform well in static environments, they struggle in dynamic environments. More recent approaches [33, 39] tackle short-term dynamics within the camera’s view by incorporating segmentation models to suppress transient objects like people or small moving elements. However, they remain limited in addressing long-term scene changes, as stale observations corrupt the optimization process. In contrast, GaME is designed to robustly handle evolving environments by filtering outdated information and maintaining high rendering quality throughout reconstruction.
@@ 3 Background: 3D Gaussian Splatting
3D Gaussian splatting (3DGS) [12] is an effective method for representing 3D scenes with novel-view synthesis capability. This approach is notable for its speed, without compromising the rendering quality. In Kerbl et al. [12], 3D Gaussians are initialized from sparse Structure-from-Motion points and are further optimized using differentiable rendering from multiple views. Each Gaussian is parameterized by mean $\mu\in\mathbb{R}^{3}$, covariance $\Sigma\in\mathbb{R}^{3\times 3}$, opacity $o\in\mathbb{R}$, and RGB color $C\in\mathbb{R}^{3}$. The mean of a projected (splatted) 3D Gaussian in the 2D image plane $\mu^{I}$ is computed as follows:
$\mu^{I}=\pi\big{(}P(T_{wc}\mu_{\text{homogeneous}})\big{)}\enspace,$ (1)
@@
$\Sigma^{I}=JR_{wc}\Sigma R_{wc}^{T}J^{T}\enspace,$ (2)
where $J\in\mathbb{R}^{2\times 3}$ is an affine transformation from Zwicker et al. [41], $R_{wc}\in SO(3)$ is the rotation component of world-to-camera transformation $T_{wc}$. We refer to Zwicker et al. [41] for further details about the projection matrices. The color $C$ along one channel $ch$ at a pixel $i$ is influenced by $m$ ordered Gaussians an rendered as:
$C^{ch}_{i}=\sum_{j\leq m}C^{ch}_{j}\cdot\alpha_{j}\cdot\prod_{k<j}(1-\alpha_{k}),\text{ with }\;\alpha_{j}=o_{j}\cdot\exp(-\sigma_{j}),\text{ and }\;\sigma_{j}=\frac{1}{2}\Delta_{j}^{T}\Sigma_{j}^{I-1}\Delta_{j}\enspace,$ (3)
@@ 4 Method
GaME builds and maintains a 3D map capable of novel view synthesis of an evolving environment, i.e., one where changes can occur during scanning but outside the view of the sensor at any time. An overview of our system is shown in Fig. 1. GaME processes depth and color images from an RGB-D sensor, using camera poses and panoptic segmentation from external estimators [2, 15]. For each keyframe, GaME triggers the Dynamic Scene Adaptation (DSA) module to incorporate new geometry and update stale regions. Simultaneously, the keyframe management system updates the state of the participating keyframes to ensure consistent multi-view optimization despite scene changes.
@@ 4.1 Online Mapping
We model our map as a set of 3D Gaussians $\{\mathbf{G}_{i}\}_{i=1}^{N}$. The goal is to create a map with object-level semantic consistency [26]. Intuitively, this reflects the prior knowledge that objects tend to move as a whole, even under partial observations of changed objects. However, directly extracting object-level submaps is challenging, as noisy observations can lead to many unnecessary object allocation and de-allocation operations [26]. This problem is exacerbated for 3DGS, where optimizing individual object clouds is difficult since the splatting mechanism (3) inherently correlates all Gaussians. Instead, we propose to build a singular 3DGS representation and extract objects on an ‘as needed’ basis during Dynamic Scene Adaptation (Sec. 4.2). During online mapping, we optimize the 3DGS parameters of the scene using a set of covisible keyframes for supervision, minimizing the loss:
@@
$L_{\text{color}}(\hat{I},I)=(1-\lambda)\cdot\frac{1}{K}\sum_{p}\lVert\hat{I}(p)-I(p)\rVert+\lambda\big{(}1-\mathrm{SSIM}(\hat{I},I)\big{)},$ (5)
@@
where $K$ is the number of rendered pixels in the rendered image or depth map, $p\in\mathbb{Z}^{2}$ denotes the pixel coordinates, and $SSIM$ is a structure similarity loss [30]. Our Keyframe Management System (Sec. 4.3) ensures that new keyframes are fully incorporated into the optimization process and that pixels observing inconsistent geometry are systematically excluded from the optimization process.
@@ 4.2 Dynamic Scene Adaptation
GaME addresses three potential scenarios: the addition, movement, or removal of an object. Note that movement can be decomposed into a removal, tracking, and re-addition step. While GaME can be readily combined with object re-localization techniques [29, 8, 25], tracking does not affect NVS and the problem can thus be simplified into two core operations: Add and Remove.
Add. In contrast to adding 3DGS in static scenes, in evolving scenes, it is essential to distinguish between adding new geometry for areas that were previously unseen and for objects that have newly appeared. We identify under-reconstructed regions in the image by low Gaussian opacity and high re-rendering depth loss. Formally, regions $\mathcal{R}_{\text{unobserved}}$ where new Gaussians are added are given by:
$\mathcal{R}_{\text{unobserved}}=\{\mathbf{p}\in\mathbb{Z}^{2}\mid L_{\text{depth}}(\mathbf{p})>\theta_{\text{depth}}\wedge\alpha(\mathbf{p})<\theta_{\text{opacity}}\},$ (7)
@@
where $L_{\text{depth}}(\mathbf{p})$ denotes the re-rendering depth loss at pixel $p\in\mathbb{Z}^{2}$, $\alpha(\mathbf{p})$ represents rendered opacity, and $\theta_{\text{depth}}\in\mathbb{R}$, and $\theta_{\text{opacity}}\in\mathbb{R}$ are the respective thresholds. The method lifts the input RGB-D frame in $\mathcal{R}_{\text{unobserved}}$ to 3D and uses the resulting points to initialize the means for new Gaussians.
DSA further determines which parts of the input correspond to newly added objects, where regions of high opacity but with an input depth significantly smaller than the rendered depth indicate previously reconstructed areas occluded by newly added geometry. Formally, the set of Gaussians that might need adaptation is defined as:
$\mathcal{G}=\{\mathbf{G}_{i}\mid\alpha(\mathbf{p})\geq\theta_{\text{opacity}}\wedge\hat{D}(\mathbf{p})>D(\mathbf{p})+\epsilon_{\text{depth}}\},$ (8)
where $\epsilon_{\text{depth}}\in\mathbb{R}$ accounts for depth measurement noise and ensures robustness against Gaussians still converging to their final values. To enforce semantic consistency, GaME then examines each object mask of the current frame. If a sufficient portion of the mask overlaps with the rendering of $\mathcal{G}$, it marks the corresponding Gaussians from $\mathcal{G}$ as needing adaptation:
$\mathcal{G}_{\text{add}}=\left\{\mathbf{G}_{i}\in\mathcal{G}\mid\frac{|\text{mask}(\mathbf{p})\cap\text{render}(\mathcal{G})|}{|\text{mask}(\mathbf{p})|}\geq\theta_{\text{mask}}\right\},$ (9)
where $\text{mask}(\mathbf{p})$ denotes the binary mask indicating the presence of the object at pixel $\mathbf{p}$, and $\theta_{\text{mask}}\in\mathbb{R}$ is the overlap threshold. This further increases robustness against noisy measurements to avoid allocating spurious Gaussians or ‘floaters’.
@@
Figure 2: Illustration of Add and Remove operations. The input disagrees with the rendered model (red). For disappearance (top), the conflicting region is projected from previous keyframes for removal (red), where semantic consistency is enforced through the object mask (blue). This allows GaME to extract complete objects even under partial observations and occlusion. When a new object appears on the scene (bottom), new Gaussians are added (red) and the area of the new object is marked as stale in previous keyframes to prevent the contamination of the optimization process.
Remove. Similarly, DSA identifies objects that have disappeared from the scene. Specifically, changed parts of the model can be identified as areas where the opacity is high (the model is well reconstructed), but the model disagrees (visually or geometrically) with the measurements:
@@
$\mathcal{G}_{\text{remove}}=\{\mathbf{G}_{i}\mid\alpha(\mathbf{p})>\theta_{\text{opacity}}\wedge L_{\text{color}}(\hat{I},I)>\theta_{\text{color}}\wedge\hat{D}(\mathbf{p})<D(\mathbf{p})-\epsilon_{\text{depth}}\},$ (10)
@@
where $\theta_{\text{color}}\in\mathbb{R}$ is the threshold. There are two notable changes compared to the Add condition. First, the sign of the depth criterion is inverted, reflecting areas where the rays of the observation would penetrate into the current model, thus indicating that the model geometry can no longer be present. This also avoids spurious detections where objects are simply occluded rather than absent. Second, we leverage the high visual fidelity of 3DGS as another conflict signal by adding a color term. In contrast to most long-term mapping methods, which rely solely on 3D information [5, 25, 13, 26, 27, 22], this allows GaME to also detect changes that are not geometrically significant, such as a sheet of paper being removed from a table. To ensure semantic consistency of removed objects, it is insufficient to rely on a single frame, as it may capture only partial views, and the object cannot be semantically detected since it is already absent. To address this, GaME renders $\mathcal{G}_{\text{remove}}$ to each covisible keyframe $KF$ and verifies whether the color and depth are consistent with the model:
$\mathcal{R}_{\text{remove}}^{\text{KF}}=\{\mathbf{p}\in\mathbb{Z}^{2}\mid L_{\text{color}}(\mathbf{p})>\theta_{\text{color}}\wedge L_{\text{depth}}(\mathbf{p})>\theta_{\text{depth}}\wedge\alpha(\mathbf{p})>\theta_{\text{opacity}}\}.$ (11)
Next, each mask in the keyframe is evaluated to determine whether a significant portion intersects with the conflicting region, marking the masked Gaussians for removal:
$\mathcal{G}^{\text{KF}}_{\text{remove}}=\left\{\mathbf{G}_{i}\mid\frac{|\text{mask}(\mathbf{p})\cap\mathcal{R}_{\text{remove}}^{\text{KF}}|}{|\text{mask}(\mathbf{p})|}\geq\theta_{\text{mask}}\right\}$ (12)
@@
Finally, the joint set, $\mathcal{G}_{\text{remove}}\cup\bigcup_{\text{KF}}\mathcal{G}_{\text{remove}}^{\text{KF}}$, is removed from the global Gaussian model. This naturally allows the extraction of individual removed objects from the global model, or deletion of the Gaussians if only an up-to-date reconstruction is desired. The DSA process is illustrated in Fig. 2.

### v2

@@ Abstract
Mapping systems with novel view synthesis (NVS) capabilities, most notably 3D Gaussian Splatting (3DGS), are widely used in computer vision, as well as in various applications, including augmented reality, robotics, and autonomous driving. However, many current approaches are limited to static scenes. While recent works have begun addressing short-term dynamics (motion within the camera’s view), long-term dynamics (the scene evolving through changes out of view) remain less explored. To overcome this limitation, we introduce a dynamic scene adaptation mechanism to update 3DGS to reflect the latest changes continuously. Since maintaining consistency remains challenging due to stale observations disrupting the reconstruction process, we further propose a novel keyframe management mechanism that discards outdated observations while preserving as much information as possible. We thoroughly evaluate Gaussian Mapping for Evolving Scenes (GaME) on both synthetic and real-world datasets, achieving a 29.7% improvement in PSNR and a $3\times$-improvement in L1 depth error over the most competitive baseline.
Figure 1: GaME is a dense mapping method capable of novel view synthesis. Given a single-camera posed RGBD input stream of an evolving environment, i.e., structural changes happening outside of the camera view but during data capture, GaME reconstructs a consistent up-to-date 3D Gaussian map that can be rendered from previously unseen viewpoints. We showcase the high-fidelity 3D Gaussian map of a real-world environment, where our method effectively detects changes and accurately integrates them into the map. Our model accurately renders fine-grained objects that were moved during the recording, highlighted in blue, yellow, and purple.
@@ 1 Introduction
Visual mapping reconstructs a consistent 3D representation of an environment from monocular or multi-view imagery. It is foundational for perception and decision-making in autonomous driving, mobile robotics, and AR/VR, where spatial understanding is required for navigation, planning, and interaction. Recent advances in visual dense reconstruction [rosinol2021kimera, murai2024_mast3rslam, schmid2024rss_khronos], have improved robustness and accuracy in complex, real-world scenes. Nevertheless, significant challenges persist in handling changing environments.
@@
Recent mapping systems have been enhanced with novel view synthesis (NVS) capabilities [murai2024monogs, keetha2024splatam], enabling them to render realistic views of scenes, allowing for detailed scene exploration and the creation of high-quality virtual environments. These approaches typically assume that the scene is static, such that optimization over multiple frames is well-conditioned. However, real-world environments are predominantly dynamic and include both short-term dynamic effects (i.e., things moving within view of the camera) as well as long-term dynamics (i.e., the scene evolving through changes outside the view of the camera; for a detailed definition see [sh-ch15-dyndef]). While some of the most recent NVS approaches address short-term dynamic objects [zheng2025wildgs, xu2024dgslam], long-term dynamics remain less explored. As a result, modern reconstruction methods with NVS capabilities struggle to capture changes as the scene evolves during data capture but outside the camera view, resulting in corrupted and erroneous reconstruction.
@@
During dense mapping of long-term dynamic scenes, systems must continuously update the 3D representation to capture the scene’s evolution and maintain reliable operation. This is especially challenging compared to classical multi-session mapping, as changes can occur at any time during data collection. Recently, initial methods have started to address this challenge. A first approach is presented in Panoptic Multi-TSDFs [schmid2022panoptic], which builds semantically consistent submaps and reasons about changes at the submap level. The following works [fu2022robust, qian2022pocd, schmid2024rss_khronos] similarly focus on object-level mapping to handle evolving scenes. However, all these methods rely on map representations that cannot easily be re-rendered, preventing their use in AR/VR or digital maps where realistic rendering is essential.
This work addresses the challenge of NVS mapping using 3DGS in evolving, long-term dynamic scenes. The key insight of our method is that two main problems prevent successful 3DGS mapping in changing environments: an outdated map and stale observations. We address these limitations by introducing a dynamic scene adaptation (DSA) mechanism that detects changes and integrates them into the 3DGS map, as well as a keyframe management system that removes stale observations while minimizing the loss of information. We make the following contributions:
@@
  • We present GaME, the first NVS-capable mapping system addressing evolving, long-term dynamic scenes.
@@
  • An efficient keyframe management strategy for accurate 3D reconstruction through long-term scene changes.
  • We thoroughly evaluate GaME on synthetic and real-world data, showing a performance increase of 325% in depth and 29% in color rendering over the most competitive baselines. We release the code open-source [^1] Released upon acceptance..
@@ 2 Related Work
3D Change Detection. The goal of change detection is to handle long-term dynamic effects, i.e., changes to the scene occurring outside the view of the sensor. Typically, this is addressed in a multi-session setting. Once a map for each session is constructed, changes can be identified through scene differencing [ambrucs2014iros_meta, fehr2017icra_tsdfchange, kim2022lt, rowell2024lista]. To achieve an object-level change understanding, this has been extended using semantic information [langer2020robust], where recent trends focus on the extraction of more specialized object features, including learned shape descriptors [wald2019rio, fu2022planesdf, rowell2024lista], neural object representations [fu2022robust, zhu2024living], and language embeddings [rashid2024lifelong]. Most related to us, lu20253dgs recently presented a 3D Gaussian Splatting-based [kerbl2023gaussian_splatting] approach by re-rendering the scene to newly collected views and using EfficientSAM [xiong2024efficientsam] for 2D change detection. However, the assumption that the scene does not change during each session and offline processing is highly limiting, as changes in evolving scenes can occur at any time during the mapping process. In contrast, our approach is designed to operate online and does not impose any limitations on when the scene changes.
@@
Online Reconstruction of Evolving Scenes. Recently, methods addressing the modeling of evolving scenes in an online fashion have emerged. A first approach is presented by schmid2022panoptic, which generates local semantically consistent submaps and incrementally reasons about changes at the submap level. fu2022robust proposes neural object descriptors to build an object-level pose graph and detect changes in the graph configuration. This has recently been extended with $SE(3)$-equivariant descriptors [fu2023neuse]. qian2022pocd present a frame-to-map-tracking approach, using a probabilistic update rule to detect object-level changes, which is further extended to a variational factor-graph approach in [qian2023pov]. A unified short- and long-term dynamic reconstruction formulation is presented in Khronos [schmid2024rss_khronos], where objects are reconstructed locally and changes detected using a library of rays. Nonetheless, these methods rely on map representations that do not support realistic rendering. In contrast, GaME provides scene reconstruction capable of real-time color and depth rendering.
@@
Dynamic Gaussian Splatting. 3D Gaussian Splatting  [kerbl2023gaussian_splatting] has revolutionized novel view synthesis (NVS) by enabling photorealistic, real-time rendering at over 100 FPS. Compared to neural radiance fields [mildenhall2021nerf], 3DGS is significantly more memory-efficient and faster to optimize. The problem of dynamic or 4D GS has attracted broad interest. A series of works [das2024neural, luiten2023dynamic, wu20244dgs, yang2023deformable3dgs] optimize a canonical set of Gaussians from the initial frame and model temporal variations through a deformation field. However, these methods are limited to short video sequences, as they cannot introduce new Gaussians after the initial frame. Another class of approaches [duan_2024_4drotorgs, kitsumata2024compact3dgs, yang2023gs4d] directly models temporal Gaussians that can exist over subsets of frames. Despite their improved flexibility, these methods still require offline optimization and multi-view input, which makes scalability a significant bottleneck for high-quality dynamic reconstruction. In contrast, our mapping operates online using only a single RGB-D camera.
@@
Online Gaussian Mapping. Most related to us, 3D Gaussian Splatting has sparked a wave of online RGB-D mapping methods [murai2024monogs, keetha2024splatam, yan2023gs, hhuang2024photoslam, yugay2023gaussianslam] that adopt 3D Gaussians as their primary scene representation. While these methods perform well in static scenes, they struggle in dynamic environments. More recent approaches [xu2024dgslam, zheng2025wildgs] tackle short-term dynamics within the camera’s view by incorporating segmentation models to suppress transient objects, such as humans or small moving elements. However, they remain limited in addressing long-term scene changes, as stale observations corrupt the optimization process. In contrast, GaME is designed to robustly handle evolving environments by filtering outdated information and maintaining high rendering quality throughout reconstruction.
@@ 3 Background: 3D Gaussian Splatting
3D Gaussian splatting (3DGS) [kerbl2023gaussian_splatting] is an effective method for representing 3D scenes with novel-view synthesis capability. This approach is notable for its speed, without compromising the rendering quality. In kerbl2023gaussian_splatting, 3D Gaussians are initialized from sparse Structure-from-Motion points and are further optimized using differentiable rendering from multiple views. Each Gaussian is parameterized by mean $\mu\in\mathbb{R}^{3}$, covariance $\Sigma\in\mathbb{R}^{3\times 3}$, opacity $o\in\mathbb{R}$, and RGB color $C\in\mathbb{R}^{3}$. The mean of a projected (splatted) 3D Gaussian in the 2D image plane $\mu^{I}$ is computed as follows:
$\mu^{I}=\pi\big(P(T_{wc}\mu_{\text{homogeneous}})\big)\kern 5.0pt,$ (1)
@@
$\Sigma^{I}=JR_{wc}\Sigma R_{wc}^{T}J^{T}\kern 5.0pt,$ (2)
where $J\in\mathbb{R}^{2\times 3}$ is an affine transformation from zwicker2001surface, $R_{wc}\in SO(3)$ is the rotation component of world-to-camera transformation $T_{wc}$. We refer to zwicker2001surface for further details about the projection matrices. The color $C$ along one channel $ch$ at a pixel $i$ is influenced by $m$ ordered Gaussians an rendered as:
$C^{ch}_{i}$ $=\sum_{j\leq m}C^{ch}_{j}\,\alpha_{j}\prod_{k<j}(1-\alpha_{k}),$
$\text{where}\quad\alpha_{j}$ $=o_{j}\,e^{-\sigma_{j}},\qquad\sigma_{j}=\tfrac{1}{2}\,\Delta_{j}^{\top}\Sigma_{j}^{-1}\Delta_{j}.$ (3)
@@ 4 Method
GaME builds and maintains a 3D map capable of novel view synthesis of a static environment undergoing structural changes happening outside the sensor’s view at any time. Our method processes posed depth and color images from a single monocular RGB-D sensor. For each keyframe, GaME triggers the Dynamic Scene Adaptation (DSA) module to incorporate changed, removed, or added geometry in the global map. Simultaneously, the keyframe management system updates the observations to ensure they reflect the scene’s up-to-date state and do not disrupt the optimization process. An overview of our approach is shown in Fig. 2.
@@ 4.1 Online Mapping
The primary goal of our method is to create a 3DGS map that is consistent with the most recent observations. Unlike previous work [schmid2022panoptic] operating on the point clouds, we chose not to model objects as separate entities, as this is prohibitively expensive in a 3DGS setup. Moreover, accurately optimizing individual objects represented as 3DGS is hard due to the nature of the splatting mechanism (3), which inherently correlates all Gaussians. Instead, we propose to build a singular representation consisting of a set of 3D Gaussians $\{\mathbf{G}_{i}\}_{i=1}^{N}$ and extract changed geometry on an ‘as needed’ basis during Dynamic Scene Adaptation (Sec. 4.2). During online mapping, we optimize the 3DGS parameters of the scene using a set of covisible keyframes for supervision, minimizing the loss:
@@
$L_{\text{color}}(\hat{I},I)$ $=(1-\lambda)\,\frac{1}{K}\sum_{p}\big\lVert\hat{I}(p)-I(p)\big\rVert$
$\quad+\,\lambda\,\big(1-\mathrm{SSIM}(\hat{I},I)\big),$ (5)
@@
where $K$ is the number of rendered pixels in the rendered image or depth map, $p\in\mathbb{Z}^{2}$ denotes the pixel coordinates, and $SSIM$ is a structure similarity loss [wang2004image].
@@ 4.2 Dynamic Scene Adaptation
In addition to reconstructing newly explored regions of the scenes, GaME addresses three potential scenarios: the addition, movement, or removal of rigid geometry outside of the camera view. While GaME can be readily combined with object re-localization techniques [wald2019rio, fu2023neuse, rowell2024lista], object tracking does not affect NVS, and the problem can thus be simplified into two core operations: Add and Remove.
Add. As the first step, DSA utilizes the current observations to detect changes in the scene where new geometry has been added. They are defined as already reconstructed Gaussians where the observed depth is closer to the camera than the rendered depth from the model:
$\mathcal{G}_{\text{add}}$ $=\Big\{\,\mathbf{G}_{i}\,\Big|\,\alpha(\mathbf{p})\geq\epsilon_{\text{opacity}}\,\wedge$
$\qquad\hat{D}(\mathbf{p})>D(\mathbf{p})+\epsilon_{\text{depth}}\,\Big\},$ (7)
@@
where $\alpha(\mathbf{p})$ represents rendered opacity at pixel $\mathbf{p}\in\mathbb{Z}^{2}$, $\epsilon_{\text{opacity}},\epsilon_{\text{depth}}\in\mathbb{R}$ are the thresholds accounting for the rendering noise. DSA resets $\mathcal{G}_{\text{add}}$ using points lifted from the RGB-D frame as initialization.
On the next step, DSA dynamically seeds new Gaussian primitives in newly explored regions to reconstruct the complete scene. Following [yugay2023gaussianslam], such regions $\mathcal{R}_{\text{new}}$ are characterized by low opacity or a high rendering depth error:
$\mathcal{R}_{\text{new}}$ $=\Big\{\,\mathbf{p}\in\mathbb{Z}^{2}\,\Big|\,\alpha(\mathbf{p})<\epsilon_{\text{opacity}}\,\vee$
$\qquad L_{\text{depth}}(\mathbf{p})>\text{median}(L_{\text{depth}}(\mathbf{p}))\,\Big\}.$ (8)
where $L_{\text{depth}}(\mathbf{p})$ denotes the depth error at pixel $p\in\mathbb{Z}^{2}$. The method lifts the input RGB-D frame in $\mathcal{R}_{\text{new}}$ into 3D space and uses it as the means for new Gaussians. After that, new Gaussians are optimized together with reinitialized $\mathcal{G}_{\text{add}}$ for a small number of iterations, as described in Sec. 4.1, to integrate the changes.
@@
Figure 3: Illustration of Add and Remove operations. The input disagrees with the rendered model (red). For disappearance (top), the conflicting region is projected to previous keyframes for removal (red), where complete object consistency is enforced through the object mask (blue). This allows GaME to extract complete objects even under partial observations and occlusion. When a new object appears on the scene (bottom), new Gaussians are added (red) and the area of the new object is marked as stale in previous keyframes to prevent the contamination of the optimization process.
Remove. On the next step, DSA detects stale Gaussians that need to be removed from the scene representation. Specifically, changed parts of the 3DGS model can be identified as those with high opacity, but disagreeing visually or geometrically with the new observations:
@@
$\mathcal{G}_{\text{remove}}$ $=\Big\{\,\mathbf{G}_{i}\,\Big|\,\alpha(\mathbf{p})\geq\epsilon_{\text{opacity}}\,\wedge$ (9)
$\quad L_{\text{color}}(\hat{I},I)>\epsilon_{\text{color}}\,\wedge\,\hat{D}(\mathbf{p})<D(\mathbf{p})-\epsilon_{\text{depth}}\,\Big\}$
@@
where $\epsilon_{\text{color}}\in\mathbb{R}$ accounts for color rendering noise. There are two notable changes compared to the Add condition. First, the sign of the depth criterion is inverted, reflecting areas where the rays of the observation would penetrate into the current model, thus indicating that the model geometry can no longer be present. This also avoids spurious detections where objects are simply occluded rather than absent. Second, we leverage the high visual fidelity of 3DGS as another conflict signal by adding a color term. In contrast to most long-term mapping methods, which rely solely on 3D information [fehr2017icra_tsdfchange, rowell2024lista, kim2022lt, schmid2022panoptic, schmid2024rss_khronos, qian2023pov], this allows GaME to also detect changes that are not geometrically significant, such as a picture removed from the wall.
Naively removing the $\mathcal{G}_{\text{remove}}$ breaks the consistency of the scene. Specifically, when only parts of the objects that need to be removed are seen in the current frame. To address this, DSA renders $\mathcal{G}_{\text{remove}}$ to each covisible keyframe $KF$ and verifies whether the color and depth are consistent with the model:
$\mathcal{R}_{\text{remove}}^{\text{KF}}$ $=\Big\{\,\mathbf{p}\in\mathbb{Z}^{2}\,\Big|\,L_{\text{color}}(\mathbf{p})<\epsilon_{\text{color}},$
$\qquad L_{\text{depth}}(\mathbf{p})<\epsilon_{\text{depth}},\;\alpha(\mathbf{p})<\epsilon_{\text{opacity}}\,\Big\}.$ (10)
$\mathcal{R}_{\text{remove}}^{\text{KF}}$ highlights the areas of previous keyframes where they observe the Gaussians that are removed. However, the detected conflicting areas might not accurately cover the objects due to occlusions and sensing noise. To address this, DSA densely segments the keyframes with SAM [kirillov2023segany], and selects all masks that are intersecting with the rendering of $\mathcal{G}_{\text{remove}}$. Then the selected masks are used to segment the Gaussians that additionally need to be removed $\mathcal{G}_{\text{remove}}^{\text{KF}}$. Our method uses FlashSplat [shen2024flashsplat] to optimally assign masks to the Gaussians, providing multiview consistency making our method robust to the noisy depth and color renderings. Notably, SAM masks may overlap and carry no semantic labels; their sole purpose is to delineate the complete silhouette of object that needs to be removed. This allows DSA to remove whole objects, even when only parts of them were visible in a current frame.
@@
Finally, the joint set, $\mathcal{G}_{\text{remove}}\cup\bigcup_{\text{KF}}\mathcal{G}_{\text{remove}}^{\text{KF}}$, is removed from the global Gaussian model, and all the Gaussians are optimized for a small number of iterations to integrate the changes. The DSA process is illustrated in Fig. 3.